OpenStudy (anonymous):

Solve 6 over x minus 4 equals 4 over x for x and determine if the solution is extraneous or not. Explain @SolomonZelman

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{6}{x-4}=\frac{4}{x} }\) like this ?

OpenStudy (anonymous):

Yeah

OpenStudy (solomonzelman):

I would suggest 1) Cross multiply (and expand both sides) 2) subtract 4x from both sides 3) divide both sides by 2

OpenStudy (solomonzelman):

You can as well: 1) Divide the numerators of both sides by 2, this will give you that the numberator on the left is 3, and numerator on the right is 2. 2) Then cross multiply. 3) subtract 2x from both sides

OpenStudy (anonymous):

Alright now determine if the solution is extraneous or not

OpenStudy (solomonzelman):

same thing in different order.

OpenStudy (solomonzelman):

What solution did you get ?

OpenStudy (anonymous):

-8

OpenStudy (solomonzelman):

yup ;)

OpenStudy (anonymous):

Thanks and later ()_()

OpenStudy (solomonzelman):

Put it in, and see if it works. \(\LARGE\color{blue}{ \frac{6}{x-4}=\frac{4}{x} }\) \(\LARGE\color{blue}{ \frac{6}{\color{red}{-8}-4}=\frac{4}{\color{red}{-8}} }\) \(\LARGE\color{blue}{ \frac{6}{-12}=\frac{4}{-8} }\) \(\LARGE\color{blue}{ -\frac{1}{2}=-\frac{1}{2} }\)

OpenStudy (solomonzelman):

so it works.

OpenStudy (solomonzelman):

\(\Large\color{ blue }{\Large {\bbox[5pt, lightcyan ,border:2px solid white ]{ \rm{Anytime!}\Huge{^✍} }}}\)